# Many-Body Localization in the Age of Classical Computing

### v2

@@ III.4 Modeling and phenomenology of MBL
where $\hat{S}^{\alpha}_{i}=\sigma^{\alpha}_{i}/2$ are spin 1/2 operators ($\alpha=x,y,z$), periodic boundary conditions are assumed ($S^{\alpha}_{L+1}\equiv S^{\alpha}_{1}$), and $J_{i}=1$ sets the energy scale. Disorder is introduced to the system via random magnetic fields, $h_{i}$, which are taken as independent random variables, each distributed uniformly in the interval $[-W,W]$, where $W$ is the disorder strength. The Jordan-Wigner transformation [301] allows to map the Hamiltonian (7) to a system of spinless fermions with annihilation operators given by $\hat{c}_{j}=(\prod_{i=1}^{j-1}\hat{S}^{z}_{i})\hat{S}^{-}_{j}$ where $\hat{S}^{-}_{j}=\hat{S}^{x}_{j}-i\hat{S}^{y}_{j}$. Under this mapping, the term $\hat{S}^{z}_{i}\hat{S}^{z}_{i+1}$ becomes a density-density interaction term $\hat{n}_{i}\hat{n}_{i+1}$ in the fermionic language up to an energy shift, where $\hat{n}_{i}=\hat{c}^{\dagger}_{j}\hat{c}_{j}$ is the fermionic number operator. For concreteness, one usually fixes $\Delta=1$, however, the non-equilibrium properties of the system in presence of disorder, $W>0$, are qualitatively the same as long as $\Delta$ is of the order of unity.
Early studies [550, 125] observed that as $W$ increases, the system starts to exhibit signatures of non-ergodicity. Those numerical results were later interpreted in terms of a transition to the MBL phase [451, 460]. Since then, various features of the strongly disordered XXZ spin-1/2 chain, in particular the slow dynamics and the low-entanglement structure of many-body eigenstates, have been interpreted as features of a stable MBL phase as a dynamical phase of matter. In the latter, the system retains memory about the details of the initial state, which may be probed by density correlation functions [561, 399]. The slow dynamics of MBL phase is associated with a logarithmic increase of entanglement entropy for initial product states [697, 47, 580, 290], as well as with dephasing of local observables [578] and with local perturbations spreading along a logarithmic light cone [200, 268, 285]. Eigenvalues of the MBL system have Poisson statistics [451], similarly to Anderson localized systems [587]. Highly excited many-body eigenstates are typically multifractal (rather than localized) in the computational basis [401], and are characterized by an area-law entanglement [579, 52], similar to ground states of local Hamiltonians [190].
The LIOMs were proposed [579, 288, 2] to describe features of systems in the MBL phase. In the absence of interactions, i.e., for $\Delta=0$, the single particle eigenstates $\phi_{\alpha}$ of (7) define a set of occupation number operators
@@
The operators $\hat{n}_{\alpha}$ form a set of mutually commuting conserved quantities which are quasilocal due to the exponential decay of the single particle eigenstate amplitude $\phi_{\alpha}(i)$ from the center of mass of state $\phi_{\alpha}$. The picture of LIOMs assumes that at strong disorder $W$ and in presence of interactions, $\Delta>0$, it is possible to construct quasilocal operators that are conserved by (7) as dressed versions of (8), which can be expressed within the following ansatz [528]
@@
When the above construction of a set of LIOMs succeeds, the system may be characterized by a complete set of conserved operators, $[\hat{H},I_{\alpha}]=0$. Conservation of LIOMs immediately implies that the long-time average of a local observable $A$ is determined not only by the average energy, $\langle E\rangle$, and its spread, $\delta E$, as in (4), but also by the average values of LIOMs in the initial state, $\braket{\psi}{I_{\alpha}}{\psi}$, consistently with the definition of MBL phase (6), implying the memory of the initial state and the absence of transport. Spectrum of $\hat{H}$ is fully determined by LIOMs, which explains the Poisson statistics of eigenvalues as well as, via the quasi-locality of $I_{\alpha}$, the area-law entanglement of eigenstates of MBL systems. The exponential decay of couplings between the LIOMs corresponding to different single particle eigenstates $\phi_{\alpha}$, $\phi_{\beta}$ explains the slow spreading of entanglement and information in the MBL phase.
@@
The construction of the quasi-local unitary transformation $\hat{U}$, which maps the original particle number operators $\hat{n}_{i}$ to the LIOMs $I_{\alpha}$, is a non-trivial task. It has been attempted for specific interacting many-body systems using perturbation theory [528] (with assumptions equivalent to those made in [51]), an iterative diagonalization method [293], or numerical approaches. All of those methods are either approximate, or allow to construct LIOMs in relatively small systems. Hence, while they provide useful information about the MBL regime in specific considered system, they do not answer the question about the fate of the system in the double limit $t\to\infty$, $L\to\infty$.
@@
Another important step in modeling of MBL, going beyond perturbation theory in the interaction, was achieved by the so-called avalanche mechanism [158], which is discussed in more details in Sec. IX. The main idea of the avalanche mechanism is to describe the interplay between a small ergodic seed, which exist with probability one in regions of small disorder, and the localized qubits. The corresponding toy model can be expressed as [158]
$\hat{H}=R\otimes\mathds{1}+g_{0}\sum_{j=1}^{L}\alpha^{j}\hat{S}^{x}_{n_{j}}\hat{S}^{x}_{j}+\sum_{j=1}^{L}h_{j}\hat{S}^{z}_{j}\;,$ (11)
@@
where the first term on the r.h.s. acts non-trivially only within the small ergodic seed, the second term describes the interaction between the spin-1/2 particle at site $j$ with a randomly selected particle within the small ergodic seed ($\alpha$ is the coupling parameter), and the third term describes the impact of random magnetic field. The actual relevance of the model in (11) for the description of the disordered spin chains, and systems in higher-dimensional lattices, is still unclear. However, a recent study [694] argued that these types of models, dubbed quantum sun models, are interesting per se, since they allow for an accurate numerical determination of the ergodicity breaking transition point that agrees to high precision with the analytical prediction based on the hybridization condition [158]. Throughout this review, we will refer in several occasions to the quantum sun models, such as the one in Eq. (11), as a reference point of the ergodicity breaking transition that is adequately described by the state-of-the-art numerical simulations on classical computers.
@@ III.5 Challenges to MBL
The interpretation in terms of a stable MBL phase was widely accepted until the year 2019. In that year, a work [690] (see also [692]) suggested that the numerical data of spectral statistics for strongly disordered spin chains can be interpreted in terms of an abrupt slowing down of dynamics with increasing disorder strength $W$, which, however, may not lead to the breakdown of ergodicity and the onset of MBL phase. Instead, [690] conjectured that in the $t\to\infty$, $L\to\infty$ limit the system remains ergodic at any disorder strength $W$, in consistence with the ETH (4). This interpretation first triggered criticism focusing on implications of the arguments of [690] to the numerical results for models in which presence of localized phase is well established, such as Anderson model in higher dimensions [593] or on random regular graphs [1]. Nevertheless, as a consequence of those works, it is now accepted that the status of MBL as of a stable dynamical phase of matter is not clear [462].
Arguments about a possible instability of the MBL phase to non-perturbative effects were later also formulated using complementary measures, such as those indicating an unbounded number entropy growth [318, 319] and the fragility of the LIOMs placed in contact with large ergodic systems [573]. All the contributions discussed above lead to an intense debate about the status of MBL phase, with various approaches leading to disparate and sometimes conflicting predictions that we aim to review in more details in the next Sections. As a working definition in studies of finite systems, which feature a certain degree of non-ergodicity but one or both of the limits $t\to\infty$, $L\to\infty$ can not be properly taken into account, it was proposed to refer to them as being in the MBL regime [427, 604].
@@ IV.1 Main open questions
1. Definition of MBL. Broadly speaking, MBL can be defined as lack of thermalization of isolated many-body system in the presence of disorder and interactions. To make the definition of MBL phase precise, one has to take the double limit of infinitely long time evolution, $t\to\infty$, and of the infinite system size $L\to\infty$, as in Eq. (6). This leads to a question about establishing a relation between $t$ and $L$ that allows one to asses in a practical way, both numerically and experimentally, whether a given system thermalizes. Another question about the definition of MBL is whether the existence of LIOMs is equivalent to the onset of MBL, as defined by (6). If LIOMs can be constructed for a given system, their presence implies that (6) is satisfied. Conversely, does it mean that all systems satisfying (6) possess LIOMs? Moreover, should the Fock-space localization also be referred to as MBL?
2. Understanding of the MBL phase. If we define the MBL phase through the breakdown of ETH (6), the central question is whether such a phase of matter actually exists. Two possible scenarios for approaching the limit $L\to\infty$ are sketched in Fig. 1. The mathematical scheme indicating robustness of the MBL phase was proposed, under certain assumptions, several years ago [292]. However, as non-experts in mathematical physics, we cannot assess the validity and mathematical rigor of this scheme. Another question that remains partially open is the demonstration of the signatures of the MBL phase and its extent in numerical computations for specific many-body systems. Such a demonstration necessarily involves extrapolating the results to the asymptotic limit $t\to\infty$ and $L\to\infty$. Devising an extrapolation scheme that would be convincing and widely accepted is still to be achieved. Additional challenges involve identifying quantities and settings well-suited for studies of the MBL phase, and also determining what features of the dynamics of disordered many-body systems are smoking-gun evidence for MBL phase rather than just an illustration of slow dynamics in the presence of strong disorder.
@@
Figure 1: Finite-size scaling of the crossover from MBL to ergodic regime. The number of lattice sites is denoted by $L$ and $g$ is a parameter of the theory (in case of the disordered XXZ spin-1/2 chain from Eq. (7), $g=1/W$). (a) A scenario that gives rise to a stable MBL phase in the thermodynamic limit $L\to\infty$, (b) a scenario in which the MBL regime at finite $L$ evolves into an ergodic phase at $L\to\infty$, in that case the MBL in the system is only a prethermal phenomenon.
@@
3. Relationship between thermalization and transport. The ETH is often associated with diffusive transport. There is, however, some evidence for other types of transport in certain systems fulfilling ETH, see Sec. VIII. Is it possible to find a system in which ETH and subdiffusion actually coexist? In such a system the smooth function $f_{A}(E,\omega)$ from Eq. (3) may have a plateau at $\omega<\omega_{\rm Th}\sim L^{-z}$ that shrinks in a subdiffusive manner, with $z>2$ and $\omega_{\rm Th}$ denotes the Thouless energy. Conversely: could there be a non-vanishing transport in a system in the MBL phase [^8] In presence of LIOMs there is no transport. Hence, this question is equivalent to the question about the relationship between MBL and LIOMs.?
@@ IV.2 Difficulties in numerical approaches to MBL
One way to numerically study MBL is to compute eigenvalues and eigenvectors of Hamiltonian $\hat{H}$ of disordered many-body system and use them to calculate a quantity $\overline{X}$ which indicates the breakdown of ergodicity in the system. For concreteness, we assume that the quantity $\overline{X}$ is normalized in such a way that it is equal to $1$ if the system follows ETH and is equal to 00 in a large MBL system. Calculating $\overline{X}$ as a function of disorder strength $W$ for system size $L$ and taking average over disorder realizations, one observes a crossover [^9] We differentiate between the notions of the crossover and phase transition. Crossover is a gradual change of the considered property of the system observed at finite system size $L$. Phase transition is a singular change of the considered property of the system in the thermodynamic limit $L\to\infty$. Phase transition can be inferred from investigation of the system at finite $L$ if the results change systematically with $L$ according to a certain scaling form. between the ergodic ($\overline{X}\to 1$) and MBL ($\overline{X}\to 0$) regimes. The goal is to understand what is the fate of this crossover in the $L\to\infty$ limit and whether it indicates a phase transition between ergodic and MBL phases in the thermodynamic limit.
The fate of the crossover observed in numerical investigations at finite system sizes is clear in a scenario presented in Fig. 2(a). There, we observe an ergodic regime at $W<W^{*}$, in which $\overline{X}$ approaches the ergodic value, $\overline{X}\to 1$, with increasing system size $L$, and an MBL regime, at $W>W^{*}$, in which $\overline{X}$ monotonously decays to 00 with the increase of $L$. The crossing point, $W^{*}$, gives a good estimate of the critical disorder strength $W_{C}$ of the ETH-MBL transition at which $\overline{X}$ changes from $1$ to 00 in the thermodynamic limit $L\to\infty$. This scenario is, however, to a good approximation only realized in certain toy models of ergodicity breaking such as the quantum sun models [390, 694].
@@
Figure 3: Observing the putative ergodic-MBL transition in time domain. (a) Average value of a local observable $A(\tau)$, darker lines correspond to larger system size $L$. At disorder strength $W_{1}$, one observes a slow decay of $A(\tau)$ which persists to longer times $\tau$ with increasing $L$; this indicates the decay $A(\tau)\to 0$ in the limit $\tau\to\infty$ and $L\to\infty$. Hence, the ETH is valid in the thermodynamic limit at $W_{1}$. When $W$ is increased to $W_{2}>W_{1}$, the decay of $A(\tau)$ slows down considerably, but the increase of $L$ is accompanied with more persistent decay of $A(\tau)$. The system is in the MBL phase if $0<A_{\infty}\equiv\lim_{L\to\infty}\lim_{t\to\infty}\bar{A}(t)$. The value of $A_{\infty}(L)$ is accessible only for small $L$ and is decreasing with $L$ at fixed $W$, see (b), where the darker lines correspond to larger $L$.
@@
Examination of time evolution of an average value of a local observable $A(\tau)$ is a different way of investigating MBL. While relevant for numerical simulations, it is especially important due to its experimental relevance. In the presence of interactions, $A(\tau)$ averaged over disorder realizations, after an initial transient, follows a non-oscillatory time dependence, as shown in Fig. 3(a). Hence, it suffices to consider $A(\tau)$, rather than the long time average $\overline{A}(t)$, c.f. Eq. (4), to assess the ergodicity of the system. Let us assume that the observable and the initial state of a disordered many-body system are chosen in such a way that $A(\tau=0)=1$ and that $\lim_{t\to\infty}\bar{A}(t)=0$ if the system is ergodic (4). Then, according to (6), a non-zero value of $A(\tau)$ in the double limit of $L\to\infty$ and $\tau\to\infty$ signifies that the system is in the MBL phase.
@@
A typical situation encountered in numerical (and in experimental) studies of disordered many-body systems is illustrated in Fig. 3(a). At small disorder strengths, $W<W_{1}$, the quantity $A(\tau)$ quickly approaches 00 consistently with the thermalization of the system. While the approach to the thermal value slows down with increasing the disorder strength, the decay of $A(\tau)$ persists to longer times with increasing the system size. Hence, it is clear that $A(\tau)$ vanishes in the double limit $L\to\infty$ and $\tau\to\infty$ at $W=W_{1}$, and that the system is in the ergodic phase. The fate of the system becomes unclear upon further increase of the disorder strength $W$. For $W_{2}>W_{1}$, at times available in numerical simulations, the average value $A(\tau)$ is significantly larger than 00. This could be interpreted in terms of an onset of MBL in the system. However, even at the larger disorder strength $W=W_{2}$, the increase of system size $L$ leads to a stronger decay of $A(\tau)$ that persists to longer times. Thus, in absence of additional information, the value of $A(\tau)$ in the $L\to\infty$, $\tau\to\infty$ limit remains obscure, preventing one from deciding whether the system is in an MBL phase at $W=W_{2}$, or whether the ergodicity is restored at sufficiently large length and time scales.
@@
In Sec. VI, we will provide concrete examples of quantities that follow the above described patterns leading to difficulties in pin-pointing the MBL transition. Investigations of MBL so far show that also quantities more complex than the quantity $\overline{X}$ share similar non-trivial trends with the increase of time and length scales. This reflects the generic feature of numerical simulations of disordered many-body systems that reveal a persistent tendency towards thermalization in the system which strengthens with the increase of time and system size. In addition, in the presence of strong disorder, the time dynamics becomes extremely slow. Thus, the question of the existence and position of the MBL phase is inevitably tied to a formidable task of distinguishing between a very slow and completely frozen dynamics. While the overall features of ergodicity breaking indicators outlined in the above description are shared by the known observables and many-body systems, the question about identification of more favourable (easier to analyze) models and observables remains open.

### v3

@@ III.4 Modeling and phenomenology of MBL
where $\hat{S}^{\alpha}_{i}=\hat{\sigma}^{\alpha}_{i}/2$ are spin 1/2 operators ($\alpha=x,y,z$), periodic boundary conditions are assumed ($\hat{S}^{\alpha}_{L+1}\equiv\hat{S}^{\alpha}_{1}$), and $J_{i}=1$ sets the energy scale. Disorder is introduced to the system via random magnetic fields, $h_{i}$, which are taken as independent random variables, each distributed uniformly in the interval $[-W,W]$, where $W$ is the disorder strength. The Jordan-Wigner transformation [327] allows to map the Hamiltonian (7) to a system of spinless fermions with annihilation operators given by $\hat{c}_{j}=(\prod_{i=1}^{j-1}\hat{S}^{z}_{i})\hat{S}^{-}_{j}$ where $\hat{S}^{-}_{j}=\hat{S}^{x}_{j}-i\hat{S}^{y}_{j}$. Under this mapping, the term $\hat{S}^{z}_{i}\hat{S}^{z}_{i+1}$ becomes a density-density interaction term $\hat{n}_{i}\hat{n}_{i+1}$ in the fermionic language up to an energy shift, where $\hat{n}_{i}=\hat{c}^{\dagger}_{j}\hat{c}_{j}$ is the fermionic number operator. For concreteness, one usually fixes $\Delta=1$, however, the non-equilibrium properties of the system in presence of disorder, $W>0$, are qualitatively the same as long as $\Delta$ is of the order of unity.
Early studies [596, 162] observed that as $W$ increases, the system starts to exhibit signatures of non-ergodicity. Those numerical results were later interpreted in terms of a transition to the MBL phase [489, 498]. Since then, various features of the strongly disordered XXZ spin-1/2 chain, in particular the slow dynamics and the low-entanglement structure of many-body eigenstates, have been interpreted as features of a stable MBL phase as a dynamical phase of matter. In the latter, the system retains memory about the details of the initial state, which may be probed by density correlation functions [608, 435]. The slow dynamics of MBL phase is associated with a logarithmic increase of entanglement entropy for initial product states [753, 50, 627, 314], as well as with dephasing of local observables [625] and with local perturbations spreading along a logarithmic light cone [210, 287, 306, 682]. Spectrum of the MBL system has Poisson statistics [489], similarly to Anderson localized systems [636]. Highly excited many-body eigenstates are typically multifractal (rather than localized) in the computational basis [437], and are characterized by an area-law entanglement [626, 55], similar to ground states of local Hamiltonians [200].
The LIOMs were proposed [626, 310, 2] to describe features of systems in the MBL phase. In the absence of interactions, i.e., for $\Delta=0$, the single particle eigenstates $\phi_{\alpha}$ of (7) define a set of occupation number operators,
@@
The operators $\hat{n}_{\alpha}$ form a set of mutually commuting conserved quantities, which are quasilocal due to the exponential decay of the single particle eigenstate amplitude $\phi_{\alpha}(i)$ from the center of mass. The picture of LIOMs assumes that at strong disorder $W$ and in presence of interactions, $\Delta>0$, it is possible to construct quasilocal operators that are conserved by (7) as dressed versions of (8), which can be expressed within the following ansatz [574],
@@
When the above construction of a set of LIOMs succeeds, the system may be characterized by a complete set of conserved operators, $[\hat{H},\hat{I}_{\alpha}]=0$. Conservation of LIOMs immediately implies that the long-time average of a local observable $A$ is determined not only by the average energy, $\langle E\rangle$, and its spread, $\delta E$, as in (4), but also by the average values of LIOMs in the initial state, $\braket{\psi}{\hat{I}_{\alpha}}{\psi}$, consistently with the definition of MBL phase (6), implying the memory of the initial state and the absence of transport. Spectrum of $\hat{H}$ is fully determined by LIOMs, which explains the Poisson statistics of eigenvalues as well as, via the quasi-locality of $\hat{I}_{\alpha}$, the area-law entanglement of eigenstates of MBL systems. The exponential decay of couplings between the LIOMs corresponding to different single particle eigenstates explains the slow spreading of entanglement and other information measures in the MBL phase.
@@
The construction of the quasi-local unitary transformation $\hat{U}$, which maps the original particle number operators $\hat{n}_{i}$ to the LIOMs $\hat{I}_{\alpha}$, is a non-trivial task. It has been attempted for specific interacting many-body systems using perturbation theory [574] (with assumptions equivalent to those made in [54]), or numerical approaches. All of those methods are either approximate, or allow to construct LIOMs in relatively small systems. Hence, while they provide useful information about the MBL regime in a specific considered system, they do not answer the question about the fate of the system in the double limit $t\to\infty$, $L\to\infty$. Of a different nature is the multiscale analysis of [316, 573], in which LIOMs are constructed by a superconvergent method a la Kolmogorov, in which a local unitary is built iteratively to order $\lambda,\lambda^{2},\lambda^{4},...$ in the dimensionless interaction strength $\lambda$. The unitary in question rotates the number operators $\hat{n}_{\alpha}$ into $\hat{I}_{\alpha}$ for which $[\hat{H},\hat{I}_{\alpha}]=0$, and one needs to prove convergence of the procedure and locality of the result. These papers have not been vetted yet by the mathematical physics community, and we therefore refrain from making strong statements based on their results.
@@
Another important step in modeling of MBL, going beyond perturbation theory in the interaction, was achieved by the so-called avalanche mechanism [165], which is discussed in more details in Sec. IX. The main idea of the avalanche mechanism is to describe the interplay between a small ergodic seed, which exists with probability one in regions of small disorder, and the localized qubits. Denoting the Hilbert space of the ergodic seed and of the localized qubits, respectively, by $\mathcal{H}_{\mathrm{seed}}$ and $\mathcal{H}_{\mathrm{loc}}$, the Hamiltonian of the corresponding toy model acting on $\mathcal{H}_{\mathrm{seed}}\otimes\mathcal{H}_{\mathrm{loc}}$ reads [165]
[…]
@@
where the first term on the r.h.s. acts non-trivially only within the small ergodic seed, the second term describes the interaction between the spin-1/2 at site $j$ with a randomly selected spin $n(j)$ within the ergodic seed ($\alpha$ is the coupling parameter and $n(j)$ for different $j$ are uncorrelated), and the third term describes the impact of random magnetic field on the localized spins. The actual relevance of the model in (11) for the description of the disordered spin chains, and systems in higher-dimensional lattices, is still unclear. However, a recent study [750] argued that these types of models, dubbed quantum sun models, are interesting per se, since they allow for an accurate numerical determination of the ergodicity breaking transition point that agrees to high precision with the analytical prediction based on the hybridization condition [165]. Throughout this review, we will refer in several occasions to the quantum sun models, such as the one in Eq. (11), as a reference point of the ergodicity breaking transition that is adequately described by the state-of-the-art numerical simulations on classical computers.
@@ III.5 Challenges to MBL
The interpretation in terms of a stable MBL phase was widely accepted until the year 2019. In that year, a paper [746] (see also [748]) suggested that the numerical results of spectral statistics for strongly disordered spin-1/2 chains can be interpreted in terms of an abrupt slowing down of dynamics with increasing disorder strength $W$, which, however, may not lead to the breakdown of ergodicity and the onset of MBL phase. Instead, [746] conjectured that in the $t\to\infty$, $L\to\infty$ limit the system remains ergodic at any disorder strength $W$, in consistence with the ETH (4). This interpretation first triggered criticism focusing on implications of the arguments of [746] to the numerical results for models in which presence of localized phase is well established, such as Anderson model in higher dimensions [642] or on random regular graphs [1]. Nevertheless, as a consequence of those works, it is now accepted that the status of MBL as of a stable dynamical phase of matter is not clear [500].
Arguments about a possible instability of the MBL phase to non-perturbative effects were later also formulated using complementary measures, such as those indicating an unbounded number entropy growth [344, 345], dynamical constraints to localization [619], and the fragility of the LIOMs placed in contact with large ergodic systems [620]. All the contributions discussed above lead to an intense debate about the status of MBL phase, with various approaches leading to disparate and sometimes conflicting predictions that we aim to review in more details in the next Sections. As a working definition in studies of finite systems, which feature a certain degree of non-ergodicity but one or both of the limits $t\to\infty$, $L\to\infty$ can not be properly taken into account, it was proposed to refer to them as being in the MBL regime [463, 653].
@@ IV.1 Main open questions
Figure 1: Finite-size scaling of the crossover from MBL to ergodic regime. The number of lattice sites is denoted by $L$ and $g$ is a parameter of the theory (in case of the disordered XXZ spin-1/2 chain from Eq. (7), $g=1/W$). (a) A scenario that gives rise to a stable MBL phase in the thermodynamic limit $L\to\infty$, (b) a scenario in which the MBL regime at finite $L$ evolves into an ergodic phase at $L\to\infty$, in that case the MBL in the system is only a prethermal phenomenon and the thermalization occurs at long times and large system sizes.
1. Definition of MBL. Broadly speaking, MBL can be defined as lack of thermalization of isolated many-body system in the presence of disorder and interactions. To make the definition of MBL phase precise, one has to take the double limit of infinitely long time evolution, $t\to\infty$, and of the infinite system size $L\to\infty$, as in Eq. (6). This leads to a question about establishing a relation between $t$ and $L$ that allows one to assess in a practical way, both numerically and experimentally, whether a given system thermalizes. Another question about the definition of MBL is whether the existence of LIOMs is equivalent to the onset of MBL, as defined by (6). If LIOMs can be constructed for a given system, their presence implies that (6) is satisfied. Conversely, does it mean that all systems satisfying (6) possess LIOMs? Moreover, what is the relation of the Fock-space localization to MBL? Is the former a stronger and distinct form of ergodicity breaking?
2. Understanding of the MBL phase. If we define the MBL phase through the breakdown of ETH (6), the central question is whether such a phase of matter actually exists. Two possible scenarios for approaching the limit $L\to\infty$ are sketched in Fig. 1. The mathematical scheme indicating robustness of the MBL phase was proposed, under certain assumptions, several years ago [316]. A recent, subsequent analysis [573], made under weaker assumptions, rules out diffusion at strong disorder in certain quantum spin-1/2 chains, but does not exclude subdiffusive transport. However, as non-experts in mathematical physics, we cannot assess the validity and mathematical rigor of schemes. Another question that remains partially open is the demonstration of the signatures of the MBL phase and its extent in numerical computations for specific many-body systems. Such a demonstration necessarily involves extrapolating the results to the asymptotic limit $t\to\infty$ and $L\to\infty$. Devising an extrapolation scheme that would be convincing and widely accepted is still to be achieved. Additional challenges involve identifying quantities and settings well-suited for studies of the MBL phase, and also determining what features of the dynamics of disordered many-body systems are smoking-gun evidence for MBL phase rather than just an illustration of slow dynamics in the presence of strong disorder.
@@
3. Relationship between thermalization and transport. The ETH is often associated with diffusive transport. There is, however, some evidence for other types of transport in certain systems fulfilling ETH, see Sec. VIII. Is it possible to find a system in which ETH and subdiffusion actually coexist? In such a system the smooth function $f_{A}(E,\omega)$ from Eq. (3) may have a plateau at $\omega<\omega_{\rm Th}\sim L^{-z}$ that shrinks in a subdiffusive manner, with $z>2$ and $\omega_{\rm Th}$ denotes the Thouless energy. Conversely: could there be a non-vanishing transport in a system in the MBL phase [^15] In presence of LIOMs there is no transport. Hence, a non-vanishing transport in the MBL phase would exclude the existence of LIOMs. ?
@@ IV.2 Difficulties in numerical approaches to MBL
One way to numerically study MBL is to compute eigenvalues and eigenvectors of Hamiltonian $\hat{H}$ of a disordered many-body system and use them to calculate a quantity $\overline{X}$ that indicates the breakdown of ergodicity in the system. For concreteness, we assume that the quantity $\overline{X}$ is normalized in such a way that it is equal to $1$ if the system follows ETH and is equal to 00 in a large MBL system. Calculating $\overline{X}$ as a function of disorder strength $W$ for system size $L$ and taking average over disorder realizations, one observes a crossover [^16] We differentiate between the notions of the crossover and phase transition. Crossover is a gradual change of the considered property of the system observed at finite system size $L$. Phase transition is a singular change of the considered property of the system in the thermodynamic limit $L\to\infty$. Phase transition can be inferred from investigation of the system at finite $L$ if the results change systematically with $L$ according to a certain scaling form. between the ergodic ($\overline{X}\to 1$) and MBL ($\overline{X}\to 0$) regimes. The goal is to understand what is the fate of this crossover in the $L\to\infty$ limit and whether it indicates a phase transition between ergodic and MBL phases in the thermodynamic limit.
@@
The fate of the crossover observed in numerical investigations at finite system sizes is clear in a scenario presented in Fig. 2(a). There, we observe an ergodic regime at $W<W^{*}$, in which $\overline{X}$ approaches the ergodic value, $\overline{X}\to 1$, with increasing system size $L$, and an MBL regime at $W>W^{*}$, in which $\overline{X}$ monotonously decays to 00 with the increase of $L$. The crossing point, $W^{*}$, gives a good estimate of the critical disorder strength $W_{C}$ of the ETH-MBL transition at which $\overline{X}$ changes from $1$ to 00 in the thermodynamic limit $L\to\infty$. This scenario is, however, to a good approximation only realized in certain toy models of ergodicity breaking such as the quantum sun models [427, 750].
@@
Figure 3: Observing the putative ergodic-MBL transition in time domain. (a) Average value $A(\tau)$ of a local observable $\hat{A}$, darker lines correspond to larger system size $L$. At disorder strength $W_{1}$, one observes a slow decay of $A(\tau)$ which persists to longer times $\tau$ with increasing $L$; this indicates the decay $A(\tau)\to 0$ in the limit $\tau\to\infty$ and $L\to\infty$. Hence, the ETH is valid in the thermodynamic limit at $W_{1}$. When $W$ is increased to $W_{2}>W_{1}$, the decay of $A(\tau)$ slows down considerably, but the increase of $L$ is accompanied with more persistent decay of $A(\tau)$. The system is in the MBL phase if $0<A_{\infty}\equiv\lim_{L\to\infty}\lim_{t\to\infty}\bar{A}(t)$. The value of $A_{\infty}(L)$ is accessible only for small $L$ and is decreasing with $L$ at fixed $W$, see (b), where the darker lines correspond to larger $L$.
@@
Examination of time evolution of an average value $A(\tau)$ of a local observable $\hat{A}$ is a different way of investigating MBL While relevant for numerical simulations, it is especially important due to its experimental relevance. In the presence of interactions, $A(\tau)$ averaged over disorder realizations, after an initial transient, follows a non-oscillatory time dependence, as shown in Fig. 3(a). Hence, it suffices to consider $A(\tau)$, rather than the long time average $\overline{A}(t)$, c.f. Eq. (4), to assess the ergodicity of the system. Let us assume that the observable and the initial state of a disordered many-body system are chosen in such a way that $A(\tau=0)=1$ and that $\lim_{t\to\infty}\bar{A}(t)=0$ if the system is ergodic (4). Then, according to (6), a non-zero value of $A(\tau)$ in the double limit of $L\to\infty$ and $\tau\to\infty$ signifies that the system is in the MBL phase.
@@
A typical situation encountered in numerical (and in experimental) studies of disordered many-body systems is illustrated in Fig. 3(a). At small disorder strengths, $W\ll W_{1}$, the quantity $A(\tau)$ quickly approaches 00 consistently with the thermalization of the system. While the approach to the thermal value slows down with increasing the disorder strength, the decay of $A(\tau)$ persists to longer times with increasing the system size. Hence, it is clear that $A(\tau)$ vanishes in the double limit $L\to\infty$ and $\tau\to\infty$ at $W=W_{1}$, and that the system is in the ergodic phase. The fate of the system becomes unclear upon further increase of the disorder strength $W$. For $W_{2}>W_{1}$, at times available in numerical simulations, the average value $A(\tau)$ is significantly larger than 00. This could be interpreted in terms of an onset of MBL in the system. However, even at the larger disorder strength $W=W_{2}$, the increase of system size $L$ leads to a stronger decay of $A(\tau)$ that persists to longer times. Thus, in absence of additional information, the value of $A(\tau)$ in the $L\to\infty$, $\tau\to\infty$ limit remains obscure, preventing one from deciding whether the system is in an MBL phase at $W=W_{2}$, or whether the ergodicity is restored at sufficiently large length and time scales.
@@
In Sec. VI, we will provide concrete examples of quantities that follow the above described patterns leading to difficulties in pin-pointing the MBL transition. Investigations of MBL so far show that also quantities more complex than the quantity $\overline{X}$ share similar non-trivial trends with the increase of time and length scales. This reflects the generic feature of numerical simulations of disordered many-body systems that reveal a persistent tendency towards thermalization in the system which strengthens with the increase of time and system size. In addition, in the presence of strong disorder, the dynamics become extremely slow. Thus, the question of the existence and position of the MBL phase is inevitably tied to a formidable task of distinguishing between very slow and completely frozen dynamics. While the overall features of ergodicity breaking indicators outlined in the above description are shared by the known observables and many-body systems, the question about identification of more favourable (easier to analyze) models and observables remains open.